A researcher is interested in developing a model that can be used to distribute assistance to​ low-income families for food costs. She used data from a national social survey to predict weekly amount spent on food using household income​ (in $1000). The resulting regression equation is ModifyingAbove Food divided by wk with caret equals 101.33 plus 0.77 HIncome.Food/wk=101.33+0.77HIncome. How much money would be needed to feed a family for a week whose household income is​ $12,000?
A.
​$193.73
B.
​$110.57
C.
​$871.33
D.
​$9341.33
Answer #1

The correct option is B. $110.57. [ANSWER]

Explanation:
We are given the following regression equation:

Food/wk (in $) = 101.33 + 0.77*Household Income(in $1000)

Now, if we want to find how much money would be needed to feed a family for a week whose household income is ​$12,000, we replace the 'Household Income' in the above equation by 12 and we get:

Food/wk (in $) = 101.33 + 0.77*12 = 101.33 + 9.24 = $ 110.57 [ANSWER]

For any queries, feel free to comment and ask.

If the solution was helpful to you, don't forget to upvote it by clicking on the 'thumbs up' button.